OpenStudy (anonymous):

A circle has a radius of 6 in. Find the area of its inscribed equilateral triangles.

OpenStudy (anonymous):

The circumscribed equilateral triangle will have an area of:

hero (hero):

If a is one of the sides of the equilateral, then: \[r = a \frac{\sqrt{3}}{6}\] If r = 6, simply substitute r into the equation and solve for a. Then use Heron's formula to find the area.

Directrix (directrix):

 Area of inscribed triangle = 1/2 ap where a is the apothem and p is the perimeter A = 1/2 (3) (6 √ 3) (3) = 27 √ 3 ------------------- Area of Circumscribed Triangle --------------------------- A = 1/2 a (p) A = 1/2 (6) (12√ 3) * 3 A = 108 √ 3 Note. The circumscribed and inscribed equilateral triangles are similar. The apothem of the inscribed triangle is 3 and the apothem of the circumscribed triangle is 6. The ratio of the areas of the triangles is (3/6)^2 - (1/2)^2 = 1/4. Because the area of the inscribed triangle is 27√ 3, the area of the circumscribed triangle is 108 √ 3.